# Robust 6DoF Pose Estimation Against Depth Noise and a Comprehensive Evaluation on a Mobile Dataset

### v4

@@ Abstract
Robust 6DoF pose estimation with mobile devices is the foundation for applications in robotics, augmented reality, and digital twin localization. In this paper, we extensively investigate the robustness of existing RGBD-based 6DoF pose estimation methods against varying levels of depth sensor noise. We highlight that existing 6DoF pose estimation methods suffer significant performance discrepancies due to depth measurement inaccuracies. In response to the robustness issue, we present a simple and effective transformer-based 6DoF pose estimation approach called DTTDNet, featuring a novel geometric feature filtering module and a Chamfer distance loss for training. Moreover, we advance the field of robust 6DoF pose estimation and introduce a new dataset – Digital Twin Tracking Dataset Mobile (DTTD-Mobile), tailored for digital twin object tracking with noisy depth data from the mobile RGBD sensor suite of the Apple iPhone 14 Pro. Extensive experiments demonstrate that DTTDNet significantly outperforms state-of-the-art methods at least 4.32, up to 60.74 points in ADD metrics on the DTTD-Mobile. More importantly, our approach exhibits superior robustness to varying levels of measurement noise, setting a new benchmark for the robustness to noise measurements.
@@ 1 Introduction
Six-degrees-of-freedom (6DoF) object pose estimation aims at determining the position and orientation of an object in 3D space. In contrast to the more matured technology of camera tracking in static settings known as visual odometry or simultaneous localization and mapping (Campos et al., 2021; Schöps et al., 2019; Permozer & Orehovački, 2019; Keselman et al., 2017), identifying the relative position and orientation of one or more objects in respect to the user’s ego position is a core function that would ensure the quality of user experience in applications like augmented reality (AR). In the most general setting, each object with respect to the ego position may undergo independent rigid-body motion, and the combined effect of overlaying multiple objects in the scene may also cause parts of the objects to be occluded from the measurement of the ego position. In this paper, the main topic of our investigation is to study the 6DoF pose estimation problem under the most general motion, occlusion, color, and lighting conditions, especially improving the accuracy and robustness of algorithms under novel data sensor properties. The dataset and proposed model are made publicly available as open-source code at: https://github.com/augcog/DTTD2.
@@
Recent advancements in the field of 6DoF pose estimation have primarily been motivated by deep neural network (DNN) approaches that advocate end-to-end training to carry out crucial tasks such as image semantic segmentation, object classification, and object pose estimation. Notable studies (Wang et al., 2019; He et al., 2020, 2021c; Jiang et al., 2022; Mo et al., 2022) have demonstrated the effectiveness of these pose estimation algorithms using established real-world 6DoF pose estimation datasets (Hinterstoisser et al., 2013; Marion et al., 2018; Xiang et al., 2018; Liu et al., 2020; Hodaň et al., 2017; Liu et al., 2021). However, it should be noted that these datasets primarily focus on robotic grasping tasks, and applying these solutions to environments served with mobile devices introduces a fresh set of challenges. A previous work (Feng et al., 2023) first studied this gap in the context of 6DoF pose estimation and replicated real-world digital-twin scenarios with varying levels of capturing distances, lighting conditions, and object occlusions. It is important to mention that, however, this dataset was collected using Microsoft Azure Kinect, which may not be the most suitable camera platform for studying 3D localization under realistic mobile environments.
Alternatively, Apple has emerged as a strong proponent of utilizing RGB-D spatial sensors for mobile AR applications with the design of their iPhone Pro camera suite, such as on the latest iPhone 14 Pro model. This particular smartphone has been given a back-facing LiDAR depth sensor (Ilci & Toth, 2020; Li et al., 2016; Bu et al., 2019; Gu et al., 2021; You et al., 2019; Weng & Kitani, 2019), a critical component to achieving accurate and detailed 3D perception and spatial understanding. However, one distinguishing drawback of the iPhone LiDAR depth is the low resolution of the depth map provided by the iPhone ARKit(Permozer & Orehovački, 2019), a $256\times 192$ resolution compared to a $1280\times 720$ depth map provided by the Microsoft Azure Kinect. This low resolution is exacerbated by large errors in the retrieved depth map. The large amounts of error in the iPhone data also pose challenges for researchers to develop a pose estimator that can correctly predict object poses that rely heavily on the observed depth map, which has not been particularly addressed in previous works (Wang et al., 2019; Sun et al., 2022; Gu et al., 2021; You et al., 2019; Weng & Kitani, 2019; Mo et al., 2022).
@@
    We conduct extensive experiments and ablation studies to demonstrate the efficacy of DTTDNet and shed light on how the depth-robustifying module works.
@@ 2 Methods
In this section, we will elaborate on the specific details of our methods. The objective is to estimate the 3D location and pose of a known object in the camera coordinates from the RGBD images. This position can be represented using homogeneous transformation matrix $p\in SE(3)$, which consists of a rotation matrix $R\in SO(3)$ and a translation matrix $t\in\mathbb{R}^{3}$, $p=[R|t]$. Section 2.1 describes our transformer-based model architecture. Section 2.2 introduces two depth robustifying modules on depth feature extractions, dedicated to geometric feature reconstruction and filtering. Section 2.3 illustrates our modality fusion design for the model to disregard significant noisy depth feature. Finally, Section 2.4 describes our final learning objective.
@@ 2.1 Architecture Overview
Fig. 2 illustrates the overall architecture of the proposed DTTDNet. The DTTDNet pipeline takes segmented depth maps and cropped RGB images as input. It then obtains feature embedding for both RGB and depth images through separate CNN and point-cloud encoders on cropped RGB images and reconstructed point cloud corresponding to the cropped depth images. [^1] We preprocessed the RGB and depth images to guarantee the pixel-level correspondence between the RGB image and the depth image. The preprocessing process is detailed in Section 3. Inspired by PSPNet (Zhao et al., 2017), the image embedding network comprises a ResNet-18 encoder, which is then followed by 4 up-sampling layers acting as the decoder. It translates an image of size $H\times W\times 3$ into a $H\times W\times d_{rgb}$ embedding space. For depth feature extraction, we take segmented depth pixels and transform them into 3D point clouds with the camera intrinsic.
The 3D point clouds are initially processed using an auto-encoder inspired by the PointNet (Qi et al., 2017). The PointNet-style encoding step aims to capture geometric representations in latent space in $\mathbb{R}^{d_{1}}$. In this context, the encoder component produces two sets of features: early-stage point-wise features in $\mathbb{R}^{N\times d_{2}}$ and global geometric features in $\mathbb{R}^{d_{3}}$. Subsequently, we add a decoder that is guided by a reference point set $P$ to generate the predicted point cloud $\hat{P}$. Features extracted from the encoder are subsequently combined with the learned representations to create a new feature sequence with a dimension of $\mathbb{R}^{N\times d_{geo}}$, where $d_{geo}=d_{1}+d_{2}+d_{3}$.
@@ 2.2 Design for Robustifying Depth Data
In this section, we will introduce two modules (Fig. 3), Chamfer Distance Loss (CDL) and Geometric Feature Filtering (GFF), that enable the point-cloud encoder in DTTDNet to handle noisy and low-resolution LiDAR data robustly.
Chamfer Distance Loss (CDL). Past methods either treated the depth information directly as image channels(Mo et al., 2022) or directly extracted features from a point cloud for information extraction(Wang et al., 2019). These methods underestimated the corruption of the depth data caused by noise and error during the data collection process. To address this, we first introduce a downstream task for point-cloud reconstruction and utilize the Chamfer distance as a loss function to assist our feature embedding in filtering out noise. The Chamfer distance loss (CDL) is widely used for denoising in 3D point clouds (Hermosilla et al., 2019; Duan et al., 2019), and it is defined as the following equation between two point clouds $P\in\mathbb{R}^{N\times 3}$ and $\hat{P}\in\mathbb{R}^{N\times 3}$:
@@
Geometric Feature Filtering (GFF). Due to the non-Gaussian noise distribution in iPhone LiDAR data (Fig. 5), which should be assumed for most depth camera data, normal estimators might either get perturbed by such noisy features or interpret wrong camera-object rotations. To deal with this sensor-level error, we advocate for the integration of a Geometric Feature Filtering (GFF) module before the modality fusion module. Our approach incorporates the Fast Fourier Transform (FFT) into the geometric feature encoding. Specifically, the GFF module includes an FFT, a subsequent single layer of MLP, and finally, an inverse-FFT. By leveraging FFT, we can transpose the input sequence of geometric signals to the frequency domain, which selects significant features from noisy input signals. After that, we obtain a more refined geometric embedding that is resilient to the non-Gaussian iPhone LiDAR noise.
Figure 3: Chamfer Distance Loss (CDL) and Geometric Feature Filtering (GFF).
@@ 2.3 Attention-based RGBD Fusion
Previous papers have emphasized the importance of modality fusion (He et al., 2021c; Wang et al., 2019) and the benefits of gathering nearest points from the point cloud (Mo et al., 2022; He et al., 2021c) in RGBD-based pose estimation tasks. While the feature extractor widens each point’s receptive field, we aim for features to interact beyond their corresponding points (Wang et al., 2019) or neighboring points (He et al., 2021c). In predicting the 6DoF pose of a cuboid based on multiple feature descriptors, our focus is on attending to various corner points, rather than solely those in close proximity to each other. To this end, inspired by recent transformer-based models used for modality fusion (Dosovitskiy et al., 2021; He et al., 2021a; Paul & Chen, 2021; Li et al., 2023b; Radford et al., 2021; Kim et al., 2021; Wang et al., 2022), we leverage the self-attention mechanism (Vaswani et al., 2023) to amplify and integrate important features while disregarding the significant LiDAR noise. Specifically, our fusion part is divided into two stages: modality fusion and point-wise fusion (Fig. 2). Both of our fusion modules consist of a standard transformer encoder with linear projection, multi-head attention and layer norm. The former module utilizes the embedding from single-modal encoders and feeds them into a transformer encoder in parallel for cross-modal fusion. The latter fusion module relies on similarity scores among points. It merges all feature embedding in a point-wise manner before feeding them into a transformer encoder. Detailed design and visual analysis of 2 fusion stages are described in our supplemental materials.
@@ 2.4 Learning Objective
Predictions with low confidence will lead to a low ADD loss, but this will be balanced by a high penalty from the second term with hyper-parameter $w$. Finally, the CDL loss, as outlined in Section 2.2, undergoes joint training throughout the training process, leading us to derive our ultimate learning objective as follows:
@@
StereoOBJ-1M(Liu et al., 2021) RGB $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 393,612 182 18 1,508,327
LINEMOD(Hinterstoisser et al., 2013) RGBD $\times$ $\checkmark$ $\checkmark$ $\times$ 18,000 15 15 15,784
YCB-Video(Xiang et al., 2018) RGBD $\times$ $\checkmark$ $\checkmark$ $\times$ 133,936 92 21 613,917
DTTD(Feng et al., 2023) RGBD $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 55,691 103 10 136,226
TOD(Liu et al., 2020) RGBD $\times$ $\checkmark$ $\times$ $\times$ 64,000 10 20 64,000
LabelFusion(Marion et al., 2018) RGBD $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 352,000 138 12 1,000,000
T-LESS(Hodaň et al., 2017) RGBD $\times$ $\times$ $\checkmark$ $\times$ 47,762 - 30 47,762
@@ 3 Dataset Description
DTTD-Mobile data contain 18 rigid objects along with their textured 3D models. The data are generated from 100 scenes, each of which features one or more of the objects in various orientations and occlusion. Following (Feng et al., 2023), our data generation pipeline is consisted of using a professional OptiTrack motion capture system that captures camera pose along the scene and using Apple’s ARKit [^2] https://developer.apple.com/documentation/arkit/ framework to capture RGB images from the iPhone camera and depth information from LiDAR scanner, as illustrated in Fig. 4. After obtaining such data, we then use the open-sourced data annotation pipeline provided by (Feng et al., 2023) to annotate ground-truth object poses. Through this pipeline, the dataset offers ground-truth labels for 3D object poses and per-pixel semantic segmentation. Additionally, it provides detailed camera specifications, pinhole camera projection matrices, and distortion coefficients. Detailed features and statistics are presented in Table 1. The fact that the DTTD-Mobile dataset includes multiple sets of geometrically similar objects, each having distinct color textures, poses challenges to existing digital-twin localization solutions. To ensure compatibility with other existing datasets, some of the collected objects partially overlap with the YCB-Video (Xiang et al., 2018) and DTTD (Feng et al., 2023) datasets. Specific details on data acquisition, benchmarking, and evaluation are presented in the supplementary materials.
@@
To further quantitatively assess the depth noise of each object from the iPhone’s LiDAR, we analyze the numerical difference between LiDAR-measured depth map, which is acquired directly from iPhone LiDAR, and reference depth map, which is derived through ground truth pose annotations. Specifically, to obtain the reference depth map, we leverage ground truth annotated object poses to render the depth projections of each object. We then apply the segmentation mask associated with each object to filter out depth readings that might be compromised due to occlusion. To measure the difference between ground truth and reference depth map, we introduce the depth-ADD metric, which calculates the average of pixel-wise L1 distance between the ground truth depth map and the reference depth map in each frame. The depth-ADD value of each object at frame $n$ is calculated as follows:
@@
where $D$ denotes the LiDAR depth map and $i$ denotes the index of pixels on it. $\mathrm{depth_{LiDAR}}_{i}$ and $\mathrm{depth_{ref}}_{i}$ represent the depth values from $D$ and the corresponding depth value from the reference depth map. The set $D$ encompasses all indices $i$ under an object’s segmentation mask where both $\mathrm{depth_{LiDAR}}_{i}$ and $\mathrm{depth_{ref}}_{i}$ yield values greater than zero. The final depth-ADD value of each object is the average of such measurement across all $N$ frames:
@@
Tab. 2 includes the average depth-ADD error in each sampled object in the second column. Greater depth-ADD values indicate increased distortions and the presence of long-tail noise in the depth data. Our analysis indicates that the mean depth-ADD across all objects is around 0.25m. It is worth noticing that the depth quality varies significantly and could potentially be affected by outliers. For example, there are three objects: black_marker, blue_marker and pink_marker exhibiting greater errors in comparison with the other objects. Detailed depth analysis is reported in the supplementary materials.

### v5

@@ Abstract
Robust 6DoF pose estimation with mobile devices is the foundation for applications in robotics, augmented reality, and digital twin localization. In this paper, we extensively investigate the robustness of existing RGBD-based 6DoF pose estimation methods against varying levels of depth sensor noise. We highlight that existing 6DoF pose estimation methods suffer significant performance discrepancies due to depth measurement inaccuracies. In response to the robustness issue, we present a simple and effective transformer-based 6DoF pose estimation approach called DTTDNet [^1] This work was previously presented as a non-archival poster at the 2024 ICML DMLR Workshop. This submission does not overlap with any archival publications., featuring a novel geometric feature filtering module and a Chamfer distance loss for training. Moreover, we advance the field of robust 6DoF pose estimation and introduce a new dataset – Digital Twin Tracking Dataset Mobile (DTTD-Mobile), tailored for digital twin object tracking with noisy depth data from the mobile RGBD sensor suite of the Apple iPhone 14 Pro. Extensive experiments demonstrate that DTTDNet significantly outperforms state-of-the-art methods at least 4.32, up to 60.74 points in ADD metrics on the DTTD-Mobile. More importantly, our approach exhibits superior robustness to varying levels of measurement noise, setting a new benchmark for robustness to measurement noise. The project page is publicly available at https://openark-berkeley.github.io/DTTDNet/.
@@ 1 Introduction
Six-degrees-of-freedom (6DoF) object pose estimation aims at determining the position and orientation of an object in 3D space. In contrast to the more matured technology of camera tracking in static settings known as visual odometry or simultaneous localization and mapping [9, 46, 43, 27], identifying the relative position and orientation of one or more objects with respect to the user’s ego position is a core function essential for ensuring a high-quality user experience in applications like augmented reality (AR). In the most general setting, each object with respect to the ego position may undergo independent rigid-body motion, and the combined effect of overlaying multiple objects in the scene may also cause parts of the objects to be occluded from the measurement of the ego position. In this paper, the main topic of our investigation is to study the 6DoF pose estimation problem under a wide range of motion, occlusion, color, and lighting conditions, especially improving the accuracy and robustness of algorithms under novel data sensor properties. The dataset and proposed model are made publicly available at https://github.com/augcog/DTTD2.
@@
Recent advancements in the field of 6DoF pose estimation have primarily been motivated by deep neural network (DNN) approaches that advocate end-to-end training to carry out crucial tasks such as image semantic segmentation, object classification, and object pose estimation. Notable studies [51, 19, 20, 26, 39] have demonstrated the effectiveness of these pose estimation algorithms using established real-world 6DoF pose estimation datasets [23, 38, 56, 35, 24, 36]. However, it should be noted that these datasets primarily focus on robotic grasping tasks, and applying these solutions to environments served with mobile devices introduces a fresh set of challenges. A previous work [12] first studied this gap in the context of 6DoF pose estimation and replicated real-world digital-twin scenarios with varying levels of capture distances, lighting conditions, and object occlusions. It is important to mention that, this dataset was collected using Microsoft Azure Kinect, which may not be the most suitable camera platform for studying 3D localization under realistic mobile environments.
Alternatively, Apple has emerged as a strong proponent of utilizing RGB-D spatial sensors for mobile AR applications with the design of their iPhone Pro camera suite, such as on the latest Apple iPhone 14 Pro model. This particular smartphone is equipped with a rear-facing LiDAR depth sensor [25, 32, 4, 13, 57, 55], a critical component to achieving accurate and detailed 3D perception and spatial understanding. However, one distinguishing drawback of the iPhone LiDAR depth is the low resolution of the depth map produced by the iPhone ARKit [43], a $256\times 192$ resolution compared to a $1280\times 720$ depth map provided by the Microsoft Azure Kinect. This low resolution is exacerbated by large errors in the retrieved depth map, which is also observed in [1, 29]. The large amounts of errors (as shown in Fig. 5) in the iPhone data also pose challenges for researchers to develop a pose estimator that can correctly predict object poses that rely heavily on the observed depth map, which has not been particularly addressed in previous works [51, 47, 13, 57, 55, 39]. We will demonstrate this in following experiments (Table 3).
@@
    We conduct extensive experiments and ablation studies to demonstrate the efficacy of DTTDNet and shed light on how the depth robustifying module works.
@@ 2 Methods
In this section, we elaborate on the details of our method. The objective is to estimate the 3D location and pose of a known object in the camera coordinates from the RGBD images. This position can be represented using homogeneous transformation matrix $p\in SE(3)$, which consists of a rotation matrix $R\in SO(3)$ and a translation matrix $t\in\mathbb{R}^{3}$, $p=[R|t]$. Section 2.1 describes our transformer-based model architecture. Section 2.2 introduces two depth robustifying modules on depth feature extractions, dedicated to geometric feature reconstruction and filtering. Section 2.3 illustrates our modality fusion design for the model to disregard significant noisy depth features. Finally, Section 2.4 describes our final learning objective.
@@ 2.1 Architecture Overview
Fig. 2 illustrates the overall architecture of the proposed DTTDNet. The DTTDNet pipeline takes segmented depth maps and cropped RGB images as input. It then obtains feature embedding for both RGB and depth images through separate CNN and point-cloud encoders on cropped RGB images and reconstructed point cloud corresponding to the cropped depth images. [^2] We preprocessed the RGB and depth images to guarantee the pixel-level correspondence between the RGB image and the depth image. The preprocessing process is detailed in Section 3. Inspired by PSPNet [59], the image embedding network comprises a ResNet-18 encoder, which is then followed by 4 up-sampling layers acting as the decoder. It translates an image of size $H\times W\times 3$ into a $H\times W\times d_{rgb}$ embedding space. For depth feature extraction, we take segmented depth pixels and transform them into 3D point clouds with the camera intrinsic.
The 3D point clouds are initially processed using an auto-encoder inspired by the PointNet [44]. The PointNet-style encoding step aims to capture geometric representations in latent space in $\mathbb{R}^{d_{1}}$. In this context, the encoder component produces two sets of features: early-stage point-wise features in $\mathbb{R}^{N\times d_{2}}$ and global geometric features in $\mathbb{R}^{d_{3}}$. Subsequently, we add a decoder that is guided by a reference point set $P$ to generate the predicted point cloud $\hat{P}$. Features extracted from the encoder are subsequently combined with the learned representations to create a new feature sequence with a dimension of $\mathbb{R}^{N\times d_{geo}}$, where $d_{geo}=d_{1}+d_{2}+d_{3}$.
@@ 2.2 Design for Robustifying Depth Data
In this section, we will introduce two modules (Fig. 3) that enable the point-cloud encoder in DTTDNet to handle noisy and low-resolution LiDAR data robustly.
Chamfer Distance Loss. Past methods either treated the depth information directly as image channels [39] or directly extracted features from a point cloud for information extraction [51]. These methods underestimated the corruption of the depth data caused by noise and error during the data collection process. To address this, we first introduce a downstream task for point-cloud reconstruction and utilize the Chamfer distance as a loss function to assist our feature embedding in filtering out noise. The Chamfer distance loss is widely used for denoising in 3D point clouds [22, 11], and it is defined as the following equation between two point clouds $P\in\mathbb{R}^{N\times 3}$ and $\hat{P}\in\mathbb{R}^{N\times 3}$:
@@
Geometric Feature Filtering. Due to the non-Gaussian noise distribution in iPhone LiDAR data (Fig. 5), which should be assumed for most depth camera data, normal estimators might either get perturbed by such noisy features or interpret wrong camera-object rotations. To deal with this sensor-level error, we advocate for the integration of a geometric feature filtering (GFF) module before the modality fusion module. Our approach incorporates the fast Fourier transform (FFT) into the geometric feature encoding. Specifically, the GFF module includes an FFT, a subsequent single layer of MLP, and finally, an inverse-FFT. By leveraging FFT, we can transpose the input sequence of geometric signals to the frequency domain, which selects significant features from noisy input signals. After that, we obtain a more refined geometric embedding that is resilient to the non-Gaussian iPhone LiDAR noise.
Figure 3: Chamfer distance loss and geometric feature filtering.
@@ 2.3 Attention-based RGBD Fusion
Previous papers have emphasized the importance of modality fusion [20, 51] and the benefits of gathering nearest points from the point cloud [39, 20] in RGBD-based pose estimation tasks. While the feature extractor widens each point’s receptive field, we aim for features to interact beyond their corresponding points [51] or neighboring points [20]. In predicting the 6DoF pose of a cuboid based on multiple feature descriptors, our focus is on attending to various corner points, rather than solely those in close proximity to each other. To this end, inspired by recent transformer-based models used for modality fusion [10, 17, 41, 34, 45, 28, 52, 14], we leverage the self-attention mechanism [50] to amplify and integrate important features while disregarding the significant LiDAR noise. Specifically, our fusion part is divided into two stages: modality fusion and point-wise fusion (Fig. 2). Both of our fusion modules consist of a standard transformer encoder with linear projection, multi-head attention and layer norm. The former module utilizes the embedding from single-modal encoders and feeds them into a transformer encoder in parallel for cross-modal fusion. The latter fusion module relies on similarity scores among points. It merges all feature embedding in a point-wise manner before feeding them into a transformer encoder. Detailed design and visual analysis of 2 fusion stages are described in our supplemental materials.
@@ 2.4 Learning Objective
Predictions with low confidence will lead to a low ADD loss, but this will be balanced by a high penalty from the second term with hyper-parameter $w$. Finally, the Chamfer distance loss, as outlined in Section 2.2, undergoes joint training throughout the training process, leading us to derive our ultimate learning objective as follows:
@@
StereoOBJ-1M [36] RGB $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 393,612 182 18 1,508,327
LINEMOD [23] RGBD $\times$ $\checkmark$ $\checkmark$ $\times$ 18,000 15 15 15,784
YCB-Video [56] RGBD $\times$ $\checkmark$ $\checkmark$ $\times$ 133,936 92 21 613,917
DTTD [12] RGBD $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 55,691 103 10 136,226
TOD [35] RGBD $\times$ $\checkmark$ $\times$ $\times$ 64,000 10 20 64,000
LabelFusion [38] RGBD $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 352,000 138 12 1,000,000
T-LESS [24] RGBD $\times$ $\times$ $\checkmark$ $\times$ 47,762 - 30 47,762
@@ 3 Dataset Description
DTTD-Mobile dataset contains 18 rigid objects along with their textured 3D models. The data are generated from 100 scenes, each of which features one or more of the objects in various orientations and occlusion. Following [12], our data generation pipeline is consists of using a professional OptiTrack motion capture system that captures camera pose along the scene and using Apple’s ARKit [^3] https://developer.apple.com/documentation/arkit/ framework to capture RGB images from the iPhone camera and depth information from LiDAR scanner, as illustrated in Fig. 4. After obtaining such data, we then use the open-sourced data annotation pipeline provided by [12] to annotate ground-truth object poses. Through this pipeline, the dataset offers ground-truth labels for 3D object poses and per-pixel semantic segmentation. Additionally, it provides detailed camera specifications, pinhole camera projection matrices, and distortion coefficients. Detailed features and statistics are presented in Table 1. The fact that the DTTD-Mobile dataset includes multiple sets of geometrically similar objects, each having distinct color textures, poses challenges to existing digital-twin localization solutions. To ensure compatibility with other existing datasets, some of the collected objects partially overlap with the YCB-Video [56] and DTTD [12] datasets. Specific details on data acquisition, benchmarking, and evaluation are provided in the appendix.
@@
To further quantitatively assess the depth noise of each object from the iPhone’s LiDAR, we analyze the numerical difference between LiDAR-measured depth map, which is acquired directly from iPhone LiDAR, and reference depth map, which is derived through ground truth pose annotations. Specifically, to obtain the reference depth map, we leverage ground truth annotated object poses to render the depth projections of each object. We then apply the segmentation mask associated with each object to filter out depth readings that might be compromised due to occlusion. To measure the difference between ground truth and reference depth map, we introduce the depth-ADD metric, which calculates the average pixel-wise L1 distance between the ground truth depth map and the reference depth map in each frame. The depth-ADD value of each object at frame $n$ is calculated as follows:
@@
where $D$ denotes the LiDAR depth map and $i$ denotes the index of pixels on it. $\mathrm{depth_{LiDAR}}_{i}$ and $\mathrm{depth_{ref}}_{i}$ represent the depth values from $D$ and the corresponding depth value from the reference depth map. The set $D$ encompasses all indices $i$ under an object’s segmentation mask where both $\mathrm{depth_{LiDAR}}_{i}$ and $\mathrm{depth_{ref}}_{i}$ yield values greater than zero. The final depth-ADD value of each object is the average of such measurements across all $N$ frames:
@@
Tab. 2 includes the average depth-ADD error in each sampled object in the second column. Greater depth-ADD values indicate increased distortions and the presence of long-tail noise in the depth data. Our analysis indicates that the mean depth-ADD across all objects is around 0.25m. It is worth noticing that the depth quality varies significantly and could be affected by outliers. For example, there are three objects: black_marker, blue_marker and pink_marker exhibiting greater errors in comparison with the other objects. Detailed depth analysis is reported in the appendix.